# Improving Transferable Targeted Attacks with Feature Tuning Mixup

### v1

@@ 5 Conclusion
In this work, we study the challenging problem of enhancing transferability in targeted adversarial attacks. Our investigation reveals that feature-level augmentation strategies show promising potential for boosting transferability effectively and efficiently. We introduce Feature Tuning Mixup (FTM), a novel method that incorporates both attack-specific feature perturbations and clean features in the feature space. To optimize the learnable perturbations efficiently, we develop a momentum-based stochastic update strategy. Comprehensive experiments on the ImageNet-compatible dataset demonstrate that FTM achieves substantial improvement over current state-of-the-art methods while maintain low computational cost.
Supplementary Material
6 Additional Experimental Results
Using DN-121 or LeViT as surrogate model. We report the targeted attack success rates when using DN-121 or LeViT as the surrogate model in Table 4. The results demonstrate that, across all black-box attack scenarios, our methods consistently outperform existing approaches, regardless of whether DN-121 or LeViT is used as the surrogate model.
Ensemble-based surrogate models. We report the targeted attack success rates using ensemble-based surrogate models in Table 5. We use two settings for the ensemble-based surrogate models: RN-50 + Inc-v3 and RN-50 + LeViT. The results show that our methods can be combined with current ensemble-based methods to further improve the attack success rates.
Visualization. Figure 7 and Figure 8 present visualizations of adversarial examples crafted by different targeted transfer-based attack methods.

### v2

@@ 5 Conclusion
In this work, we study the challenging problem of enhancing transferability in targeted adversarial attacks. Our investigation reveals that feature-level augmentation strategies show promising potential for boosting transferability effectively and efficiently. We introduce Feature Tuning Mixup (FTM), a novel method that incorporates both attack-specific feature perturbations and clean features in the feature space. To optimize the learnable perturbations efficiently, we develop a momentum-based stochastic update strategy. Comprehensive experiments on the ImageNet-compatible dataset demonstrate that FTM achieves substantial improvement over current state-of-the-art methods while maintaining low computational cost.
Acknowledgments. This work was supported in part by HK RGC GRF under Grant PolyU 15201323.
Supplementary Material for
Improving Transferable Targeted Attacks with Feature Tuning Mixup
In the supplementary material, we provide additional experimental results on parameter $p$ in Chapter A.1. Then, in Chapters A.2A.3, and A.4, we present additional results demonstrating the effectiveness and efficiency of FTM across different surrogate and targeted models. Finally, Chapter A.5 includes visualizations of adversarial examples generated by our FTM.
A Additional Experiments
A.1 Supplementary results on the analysis of $p$
We have shown in Section 4.3 (Ablation Study) that the attack success rate when using clean feature mixup drops to near 00 in Figure 6 due to the low clean accuracy of the perturbed surrogate model. Table 4 shows the clean accuracy of the perturbed surrogate model with varying values of $p$. The results demonstrate that increasing $p$ leads to extremely low clean accuracy when using clean feature mixup in FTM, which explains the rapid decline of the blue line in Figure 6. In contrast, without clean feature mixup, the clean accuracy remains stable, maintaining 65.3% even at $p=1.0$. Consequently, the corresponding orange line in Figure 6 exhibits relatively stable behavior. Notably, as shown in the left plot of Figure 6 our FTM performs stably around the optimal hyperparameter $p$ and $\alpha_{max}$.
Ablation Parameter $p$
0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1.0
w/ clean feature mixup 51.9 28.3 14.0 10.8 8.2 6.0 5.4 4.9 4.3 3.8
w/o clean feature mixup 79.8 73.3 69.7 68.6 67.0 66.4 66.0 65.8 65.4 65.3
Table 4: Accuracy (%) of clean images on the perturbed surrogate model with varying parameter $p$.
A.2 Attacking black-box multimodal LLMs
FTM can be used to attack multimodal LLMs. We use four commercial LLMs for evaluation, including Qwen2-VL, Llama-3.2, Claude-3.5, and GPT-4o. We randomly select 100 images from the ImageNet-compatible dataset and generate targeted adversarial examples on ViT using RDI-FTM-E. For each generated targeted adversarial example, we use the prompt “Is this image a photo of {target label}? Yes or No?” to obtain the predictions of the LLMs. Table 5 shows that our method achieves an average attack success rate of 40.5%.
Response Qwen2-VL Llama-3.2 Claude-3.5 GPT-4o Avg
Total 100 100 100 100 100 %
Refuse to Answer 0 10 0 0 2.50 %
Uncertain 1 5 1 0 1.75 %
Attack Failed 52 42 54 73 55.25 %
Attack Succeeded 47 43 45 27 40.50 %
Table 5: Evaluation on multimodal LLMs. The targeted adversarial examples are generated on ViT using RDI-FTM-E.
A.3 Evaluation with different surrogate models
We report the targeted attack success rates when using DN-121 or LeViT as the surrogate model in Table 6. The results demonstrate that, across all black-box attack scenarios, our methods consistently outperform existing approaches, regardless of whether DN-121 or LeViT is used as the surrogate model.
We report the targeted attack success rates using ensemble-based surrogate models in Table 7. We use two settings for the ensemble-based surrogate models: RN-50 + Inc-v3 and RN-50 + LeViT. The results show that our methods can be combined with current ensemble-based methods to further improve the attack success rates.
A.4 Efficiency of FTM on attacking ViT
[…]
A.5 Visualization of targetd adversarial examples
Figure 7 and Figure 8 present visualizations of adversarial examples crafted by different targeted transfer-based attack methods.
@@ A.5 Visualization of targetd adversarial examples
Source Attack Time (s) VGG-16 RN-18 RN-50 DN-121 Xcep MB-v2 EF-B0 IR-v2 Inc-v3 Inc-v4 Avg.
ViT RDI 1.76 2.4 2.4 3.1 5.2 3.5 3.6 10.6 4.4 5.3 4.3 4.5
RDI-VT 10.20 2.5 3.6 4.1 7.1 5.9 4.1 13.4 7.4 5.5 5.4 5.9
RDI-Admix 5.11 5.4 5.4 6.9 11.5 6.7 6.9 18.7 8.1 10.1 8.0 8.8
RDI-SI 8.38 4.2 8.2 9.5 17.6 10.2 9.1 24.7 13.3 18.0 12.1 12.7
RDI-ODI 4.53 10.2 12.7 14.3 22.1 18.6 12.1 32.2 22.5 22.3 20.7 18.8
RDI-CFM 1.80 10.8 14.2 14.9 21.0 14.9 15.0 31.9 15.0 18.9 17.1 17.4
RDI-FTM 1.92 11.6 15.1 14.6 22.1 16.7 18.6 31.8 15.8 20.5 18.0 18.5
RDI-FTM-E 3.81 13.4 18.1 18.3 26.5 19.6 19.5 37.2 19.8 23.0 21.3 21.7
[…]
Source Attack Time (s) ViT LeViT ConViT Twins PiT Avg.
ViT RDI 1.76 99.5 24.3 23.3 11.7 26.0 37.0
RDI-VT 10.20 98.3 32.0 29.7 15.3 35.7 42.2
RDI-Admix 5.11 99.1 39.0 34.6 18.5 40.5 46.3
RDI-SI 8.38 98.4 58.4 69.2 38.8 63.8 65.7
RDI-ODI 4.53 87.6 45.9 42.1 29.5 49.3 50.9
RDI-CFM 1.80 96.5 47.2 50.1 27.4 48.8 54.0
RDI-FTM 1.92 94.3 50.5 53.3 31.0 52.6 56.3
RDI-FTM-E 3.81 96.0 54.9 55.6 36.3 55.8 59.7
[…]